1. A battery is connected to a light bulb, lighting the bulb. Where in the circuit is the current the greatest?

A) inside the battery
B) it's the same at all these points
C) inside the light bulb filament
D) between the light bulb and the negative terminal of the battery
E) between the positive terminal of the battery and the light bulb

2. The emf of a battery is:

A) its internal resistance
B) the voltage difference it provides
C) the power it provides
D) the current it provides
E) the force it provides

3. A wire A carrying current splits into wires B and C at a junction (the currents in A was into the junction; the currents in B and C are out of the junction). We can say that:

A) the current in A = current in B - current in C
B) the current in A = current in B + current in C
C) the current in A = current in B = current in C
D) the current in A < current in B + current in C
E) the current in A > current in B + current in C

4. A light bulb is connected to an ideal battery, so that the light bulb is lit. What can we say about the voltage across the two?

A) The voltage difference across the battery is smaller than the voltage difference across the bulb
B) The voltage difference across the battery is larger than the voltage difference across the bulb
C) The voltage difference across the battery is equal to the voltage difference across the bulb

5. You connect a length of metal across the terminals of a battery and a current i flows through it. If you were to connect two of these lengths of the metal (end-to-end) to the battery terminals, a current of ____ would flow.

A) 2i
B) i/4
C) i/2
D) i
E) 4i

6. You connect an ideal battery to a variable resistor (its resistance can by changed by, say, a knob). As you vary the resistance, what if anything in the circuit doesn't change?

A) the voltage across the resistor
B) the current through the resistor
C) the current through the battery

7. Current goes to the right through a resistor R. What can you say about the potential V and electric field E?

A) There is no E field in the resistor, V is larger on the left side of R.
B) There is an E field to the right, V is larger on the right side of R.
C) There is no E field in the resistor, V is the same on both sides of R.
D) There is no E field in the resistor, V is larger on the right side of R.

8. We can characterize a circuit element by its I-V curve (a graph of current through it vs voltage difference across it). For an Ohmic resistor, the resistance is the:

A) peak value of this graph
B) y-intercept of this graph
C) 1/slope of this graph
D) slope of this graph
E) x-intercept of this graph
E) There is an E field to the left, V is larger on the left side of R.
F) There is an E field to the left, V is larger on the right side of R.
G) There is an E field to the right, V is the same on both sides of R.
H) There is an E field to the right, V is larger on the left side of R.
Answer #1

1)

current I = E/R

A) inside the battery

2)

the voltage difference it provides

3)

from kirchoff I law

sum of all current meeting at a junction = 0


iA = iB + iC


B) the current in A = current in B + current in C

___________________

4)

dV = E-ir

for ideal battery r = 0

dV = E

C) The voltage difference across the battery is equal to the voltage difference across the bulb

_________________________


resistance = rho*L/A


L increases R increases

i = V/R


current decreases


I = i/2

option (C)


_______________


6)


from ohms law

v = iR

as R changes i changes and voltage reamins same


option A

+_____________________


7)


current flows from high to low


curre t flows in the direction E


H) There is an E field to the right, V is larger on the left side of R.

_______________


8)


V = IR

compare with y = mx


R = m = slope

D) slope of this graph